# Decomposition-Based Optimal Bounds for Privacy Amplification via Shuffling

### v1

@@ Appendix B Proof of Theorem 5
The proof of this theorem follows similar lines as the proof of Lemma 12 in (Balle et al., 2019).
Define random variables $Y_{1}^{b}\sim\mathcal{R}(x_{1}^{b})$ and $W_{i}\sim\mathcal{R}(x_{2}),i=1,2,\dots,n-1$. $\mathbf{W}_{n-1}=\{W_{1},W_{2},\dots,W_{n-1}\}$. Let $\vec{y}\in\mathbb{Y}^{n}$ be a tuple of elements from $\mathbb{Y}$ and $Y\in\mathbb{N}_{n}^{\mathbb{Y}}$ be the corresponding multiset of entries. Then we have
@@
$\mathbb{P}\left[(Y_{1}^{b},W_{1},\ldots,W_{n-1})=\vec{y}_{\sigma}\right]=\mathcal{R}(x_{1}^{b})(y_{\sigma(1)})\prod_{i=2}^{n}\mathcal{R}(x_{2})(y_{\sigma(i)}).$
Summing this expression over all permutations $\sigma$ and factoring out the product of the $\mathcal{R}(x_{2})$’s yields:
$\frac{1}{n!}\sum_{\sigma}\mathcal{R}(x_{1}^{b})(y_{\sigma(1)})\mathcal{R}(x_{2})(y_{\sigma(2)})\cdots\mathcal{R}(x_{2})(y_{\sigma(n)})$
$=\left(\prod_{i=1}^{n}\mathcal{R}(x_{2})(y_{i})\right)\left(\frac{1}{n}\sum_{i=1}^{n}\frac{\mathcal{R}(x_{1}^{b})(y_{i})}{\mathcal{R}(x_{2})(y_{i})}\right)$
$=\mathbb{P}[\mathbf{W}_{n}=Y]\cdot\frac{1}{n}\sum_{i=1}^{n}\frac{\mathcal{R}(x_{1}^{b})(y_{i})}{\mathcal{R}(x_{2})(y_{i})}.$
@@
$\mathcal{D}_{e^{\varepsilon}}\big{(}\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big{)}=\mathcal{D}_{e^{\varepsilon}}(\{Y_{1}^{0}\}\cup W_{n-1}\parallel\{Y_{1}^{1}\}\cup W_{n-1})$
$=\int_{\mathbb{N}_{n}^{\mathbb{Y}}}\left[\mathbb{P}[\{Y_{1}^{0}\}\cup W_{n-1}=Y]-e^{\varepsilon}\mathbb{P}[\{Y_{1}^{1}\}\cup W_{n-1}=Y]\right]_{+}$
$=\int_{\mathbb{N}_{n}^{\mathbb{Y}}}\mathbb{P}[W_{n}=Y]\left[\frac{1}{n}\sum_{i=1}^{n}\frac{\mathcal{R}(x_{1}^{0})(y_{i})-e^{\varepsilon}\mathcal{R}(x_{1}^{1})(y_{i})}{\mathcal{R}(x_{2})(y_{i})}\right]_{+}$
$=\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}\frac{\mathcal{R}(x_{1}^{0})(y_{i})-e^{\varepsilon}\mathcal{R}(x_{1}^{1})(y_{i})}{\mathcal{R}(x_{2})(y_{i})}\right]_{+}$
$=\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}G_{i}\right]_{+}.\qed$

### v5

@@ Appendix C Proof of Theorem 12
The proof of this theorem follows similar lines as the proof of Lemma 12 in [BBGN19].
Let the support of $A_{2}$ be $\mathbb{Y}$. Since $A_{0}$ and $A_{1}$ are absolutely continuous with respect to $A_{2}$, their supports are subsets of $\mathbb{Y}$. Define random variables $Y_{1}^{b}\sim A_{b},b=0,1$ and $W_{i}\sim A_{2},i=1,2,\dots,n-1$. $\mathbf{W}_{n-1}=\{W_{1},W_{2},\dots,W_{n-1}\}$. Let $\vec{y}\in\mathbb{Y}^{n}$ be a tuple of elements from $\mathbb{Y}$ and $Y\in\mathbb{N}_{n}^{\mathbb{Y}}$ be the corresponding multiset of entries. Then we have
@@
$\mathbb{P}\left[(Y_{1}^{b},W_{1},\ldots,W_{n-1})=\vec{y}_{\sigma}\right]=A_{b}(y_{\sigma(1)})\prod_{i=2}^{n}A_{2}(y_{\sigma(i)}),\quad b=0,1.$
Summing this expression over all permutations $\sigma$ and factoring out the product of the $A_{2}$’s yields:
$\frac{1}{n!}\sum_{\sigma}A_{b}(y_{\sigma(1)})A_{2}(y_{\sigma(2)})\cdots A_{2}(y_{\sigma(n)})$ $=\left(\prod_{i=1}^{n}A_{2}(y_{i})\right)\left(\frac{1}{n}\sum_{i=1}^{n}\frac{A_{b}(y_{i})}{A_{2}(y_{i})}\right)$
$=\mathbb{P}[\mathbf{W}_{n}=Y]\cdot\frac{1}{n}\sum_{i=1}^{n}\frac{A_{b}(y_{i})}{A_{2}(y_{i})}.$
@@
$D_{e^{\varepsilon}}\big{(}P_{0}\parallel P_{1}\big{)}$ $=D_{e^{\varepsilon}}(\{Y_{1}^{0}\}\cup\mathbf{W}_{n-1}\parallel\{Y_{1}^{1}\}\cup\mathbf{W}_{n-1})$
$=\int_{\mathbb{N}_{n}^{\mathbb{Y}}}\left[\mathbb{P}[\{Y_{1}^{0}\}\cup\mathbf{W}_{n-1}=Y]-e^{\varepsilon}\mathbb{P}[\{Y_{1}^{1}\}\cup\mathbf{W}_{n-1}=Y]\right]_{+}$
$=\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}\frac{A_{0}(y_{i})-e^{\varepsilon}A_{1}(y_{i})}{A_{2}(y_{i})}\right]_{+}$
$=\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}G_{i}\right]_{+}.$
10-RR, low $\varepsilon_{0}$
10-RR, large $\varepsilon_{0}$
BLH, low $\varepsilon_{0}$
BLH, large $\varepsilon_{0}$
RAPPOR, low $\varepsilon_{0}$
RAPPOR, large $\varepsilon_{0}$
Figure 2: Experimental Results: RR, BLH and RAPPOR
OUE, low $\varepsilon_{0}$
OUE, large $\varepsilon_{0}$
Laplace, low $\varepsilon_{0}$
Laplace, large $\varepsilon_{0}$
Figure 3: Experimental Results: OUE and Laplace
$\frac{1}{2}$+$\frac{1}{2}$, low $\varepsilon_{0}$
$\frac{1}{2}$+$\frac{1}{2}$, large $\varepsilon_{0}$
Figure 4: Experimental Results: Parallel Composition
Figure 5: Upper Bounds for RAPPOR with Varying Domain Size $D$
Figure 6: 10-RR under Poisson Subsampling (with sampling rate $r$) and Shuffling
Figure 7: Upper Bounds for Joint Composition under the adjacency relation based on Hamming distance $d$
Figure 8: Cumulative distribution of $L^{0,1}_{\varepsilon}$ for the Laplace mechanism with $\varepsilon_{0}=1.0$ and $\varepsilon=0.1$